OpenStudy (anonymous):

When 27 times x squared times z all over negative 3 times x squared times z to the fourth power is completely simplified, the exponent on the variable z is _____. Numerical Answers please?

OpenStudy (austinl):

\(\dfrac{27\times x^2\times z}{-3\times x^2\times z^4}\Rightarrow\dfrac{-9\times\cancel{x^2}\times\cancel{z}}{\cancel{-3}\times\cancel{x^2}\times z^3}\Rightarrow \dfrac{-9}{z^3}\Rightarrow -9z^{-3}\)

OpenStudy (anonymous):

Thank's.

OpenStudy (austinl):

No problem, did what I did make sense?

OpenStudy (anonymous):

Yes, You made it very clear. and Now I understand.

OpenStudy (austinl):

Very cool, good luck with any further questions. If you need help, feel free to ask.

OpenStudy (anonymous):

Thank you, and I will.

OpenStudy (anonymous):

The -3 exponent is soooo wrong.

OpenStudy (anonymous):

BUTTER!!!

OpenStudy (tjhub00):

I just got a 100% on this and the REAL answer is 3